OpenStudy (anonymous):

9 students volunteer for a committee. How many different 7-person committees can be chosen?

OpenStudy (anonymous):

Any options?

OpenStudy (anonymous):

A.181,440 B.36 C.362,880 D.1

OpenStudy (wolf1728):

You want to choose 7 people from a set of 9

OpenStudy (anonymous):

9 C 7

OpenStudy (anonymous):

@Nasa7

OpenStudy (anonymous):

You use the formula of nCr. n is the total number you are choosing from. r is the number you want to choose.

OpenStudy (wolf1728):

The formula is n! / (r! * (n-r)!) where n is 9 and r is 7

OpenStudy (anonymous):

yep @wolf1728 !

OpenStudy (anonymous):

9!/(7!*(9-7)!)

OpenStudy (wolf1728):

9! / (7! * 2!) which equals 9 * 8 / 2 = 72 / 2 = 36 which is B

OpenStudy (anonymous):

Thanks a lot the problem was that i needed a calculator @wolf1728

OpenStudy (wolf1728):

Actually, if you cancel the factors correctly you don't need a calculator

OpenStudy (anonymous):

thanks anyway , i have not took math and i forgot all of this , this is for a different class (CCP) @wolf1728

OpenStudy (anonymous):

Sorry I just saw your tag back I see @wolf1728 is helping you now :D